Two students are analyzing a compound containing copper and oxygen only. Fred’s sample weighs 3.12g and contains 2.50g of copper. Wilma’s sample weighs 1.62g and contained 1.44g of copper. Are the two students analyzing the same compound? Justify your answer.
1
Expert's answer
2020-05-24T03:10:51-0400

The formula of the analyzed samples could be presented as follows

Cux1Oy1(Fred) and Cux2Oy2 (Wilma)

Calculations of the elements quantity in Fred's sample

n(Cu) = x1 = m(Cu)/M(Cu) = 2.50 g/64 g/mol = 0.039 mol

n(O) = y1 = m(O)/M(O) = (3.12 - 2.50) g/ 16 g/mol = 0.039 mol

Elements ratio in the sample is x1/y1 = 1/1 -> the formula of the analyzed substance is CuO


Calculations of the elements quantity in Wilma's sample

n(Cu) = x2 = m(Cu)/M(Cu) = 1.44 g/64 g/mol = 0.0225 mol

n(O) = y2 = m(O)/M(O) = (1.62 - 1.44) g/ 16 g/mol = 0.01125 mol

Elements ratio in the sample is x2/y2 = 2/1 -> the formula of the analyzed substance is Cu2O


The answer is "No"